# Transversal Hamilton cycles in digraph collections

### v1

@@ 4 Stable case
In this section, by combining the results of the previous sections, we use the regularity-blow-up method in a digraph collection $\mathcal{D}$ to find a transversal directed Hamilton cycle when $\mathcal{D}$ is strongly stable or weakly stable. Let $10<\gamma,\alpha,\epsilon,\delta<$ 1. We say that $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly table. We first show that the reduced digraph collection inherits stability.
@@
Let $0<\frac{1}{n}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu,\alpha\ll\gamma,\epsilon\ll\delta\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. Let $\mathcal{R}=\mathcal{R}(\epsilon_{0},d,L_{0})$ be the reduced digraph collection of $\mathcal{G}$. If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2},\epsilon,\frac{\delta}{2})$-stable.
@@
We may assume that $\frac{1}{n}\leq\frac{1}{n_{0}}$ where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is the constant from Lemma 2.2. Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $[n]$ and a digraph collection $\mathcal{D}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq n\leq Mm+\epsilon_{0}n$ and $Mm\leq n\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Given $X\subset[L]$, we write $\hat{X}:=\bigcup_{j\in X}V_{j}$. Hence $|\hat{X}|=m|X|$.
@@
Let $d\ll\alpha^{\prime}$ and $i\in[L]$. Assume that there are sets $A,B\subset[L]$ of size at least $(\frac{1}{2}-\alpha^{\prime})L$ such that $e_{R_{i}}(A,B)\leq\alpha^{\prime}L^{2}$. Then $e_{D_{c}}(\hat{A},\hat{B})\leq\sqrt{\alpha^{\prime}}n^{2}$ for all but at most $2\sqrt{\alpha^{\prime}}m$ colors $c\in\mathcal{C}_{i}$.
@@
Let $t$ be the number of colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A},\hat{B})>\sqrt{\alpha^{\prime}}n^{2}$. Since $e(D_{c}^{\prime})-e(D_{c})\leq(3d+\epsilon_{0})n^{2}$ for all $c\in[n]$, we have
$\sqrt{\alpha^{\prime}}n^{2}t$ $\leq\sum_{c\in\mathcal{C}_{i}}e_{D_{c}}(\hat{A},\hat{B})\leq\sum_{c\in\mathcal{C}_{i}}\left(e_{D_{c}^{\prime}}(\hat{A},\hat{B})+(3d+\epsilon_{0})n^{2}\right)$
$\leq|\mathcal{C}_{i}|e_{R_{i}}(A,B)m^{2}+|\mathcal{C}_{i}|(3d+\epsilon_{0})n^{2}$
$\leq\alpha^{\prime}L^{2}m^{3}+4dmn^{2}\leq 2\alpha^{\prime}mn^{2}.$
@@
We prove that $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Suppose that there exists a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for all $i\in I$. By Definition 3.2, for each $i\in I$, there are two sets $A^{i},B^{i}\subseteq[L]$ of size at least $(\frac{1}{2}-\alpha^{2})L$ such that $e_{R_{i}}(A^{i},B^{i})<\alpha^{2}L^{2}$. Applying Claim 4.2 with $\alpha^{\prime}:=\alpha^{2}$, we get that there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i_{0}}$ such that $e_{D_{c}}(\hat{A^{i}},\hat{B^{i}})\leq\alpha n^{2}$. Notice that $|\hat{A^{i}}|,|\hat{B^{i}}|\geq(\frac{1}{2}-\alpha^{2})Lm>(\frac{1}{2}-\alpha)n$. Then there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i}$ such that $D_{c}$ is not $\alpha$-nice. Thus the number of colors $c\in[n]$ for which $D_{c}$ is not $\alpha$-nice is at least $(1-2\alpha)m|I|>(1-\gamma)n$ since $\alpha\ll\gamma$. This contradicts the fact that $\mathcal{D}$ is $(\gamma,\alpha)$-strong stable.
@@
Suppose that $\mathcal{R}$ is not $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Then there is subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$. It suffices to show that $\mathcal{R}$ is $(\epsilon,\frac{\delta}{2})$-weakly stable. Choose $i\in I$, then $R_{i}$ is $\alpha^{2}$-extremal. Based on Lemma 2.4 (i), one has $d_{R_{i}}^{+}(j),d_{R_{i}}^{-}(j)\geq(\frac{1}{2}-2\mu)L$ for all but at most $d^{\frac{1}{4}}L$ vertices $j\in[L]$. Together with Lemma 3.3, we obtain that
[…]
Note that $|\hat{Z}_{i}^{\prime}|,|\hat{Y}_{i}^{\prime}|=(\frac{1}{2}-\alpha^{2})Lm\geq(\frac{1}{2}-2\alpha)n$. In view of Claim 4.2, we know that there is $\mathcal{B}_{i}\subseteq\mathcal{C}_{i}$ with $|\mathcal{C}_{i}\setminus\mathcal{B}_{i}|\leq 2\alpha m$ such that for all $c\in\mathcal{B}_{i}$, we have $e_{D_{c}}(\hat{Z}_{i}^{\prime},\hat{Y}_{i}^{\prime})\leq\alpha n^{2}$. It follows that $D_{c}$ is $2\alpha$-extremal for each $c\in\mathcal{B}_{i}$.
[…]
Choose $c\in\mathcal{B}_{i},c^{\prime}\in\mathcal{B}_{j}$ for some $i,j\in[L]$. Suppose that $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing. We must have $i\neq j$.
[…]
$|A_{i}^{\prime}\Delta A_{j}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{A}_{j}^{\prime}|\geq|A_{c}\Delta A_{c^{\prime}}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{A}_{j}^{\prime}\Delta A_{c^{\prime}}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
@@
[…]
$|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|{\hat{W_{i}^{3}}^{\prime}}\Delta{\hat{W_{j}^{3}}^{\prime}}|\geq|W_{c}^{3}\Delta W_{c^{\prime}}^{3}|-|{\hat{W_{i}^{3}}^{\prime}}\Delta W_{c}^{3}|-|{\hat{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-21\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
and similarly, $|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{1}}^{\prime}|\geq\frac{\delta L}{2}$. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
[…]
$|A_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{{W_{j}^{3}}^{\prime}}|\geq|A_{c}\Delta W_{c^{\prime}}^{3}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
Similarly, $|B_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
@@
Let $0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. If $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
Choose additional parameters $n_{0},L_{0},\epsilon_{0},d,\theta,\beta,\lambda,\zeta$, where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is obtained from Lemma 2.2, and $\zeta$ is obtained from Lemma 3.3, such that
$1<\frac{1}{n}\ll\frac{1}{n_{0}}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu\ll\beta\ll\lambda\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1,$
and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$, where the previous lemmas hold with suitable parameters.
@@
[…]
Let $\{c_{i},c_{i}^{\prime}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a collection of distinct colors. For each $i\in[2\beta n]$, define $\mathcal{F}_{i}$ to be a directed $2$-graph with vertex set $V$ and edge set $\{(v_{1},v_{2})\in V^{2}:(c_{i},c_{i}^{\prime},v_{1},v_{2})\ \text{is totally absorbable}\}$ and $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$. Define the collection of directed multi-graphs:
@@
For every $S:=S(u_{1},u_{2},c)\in\mathbf{Z}$ and every $i\in[2\beta n]$, we have $|E(S)\cap E(\mathcal{F}_{i})|\geq 2^{-4}n^{2}$. Applying Lemma 3.1 with $t=2\beta n$ and $\epsilon=2^{-4}$, we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-9}\beta n$ for all $S\in\mathbf{Z}$. ∎

### v2

@@ 4 Stable case
In this section, we employ the regularity-blow-up method for digraph collections $\mathcal{D}$ to prove the existence of a transversal directed Hamilton cycle when $\mathcal{D}$ is either strongly or weakly stable, combining the results from previous sections. Let $0<\gamma,\alpha,\epsilon,\delta<1$. We say that $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly stable. We first show that the reduced digraph collection inherits stability.
@@
Let ${0<}\frac{1}{n}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu,\alpha\ll\gamma,\epsilon\leq\delta^{2}\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. Let $\mathcal{R}=\mathcal{R}(\epsilon_{0},d,L_{0})$ be the reduced digraph collection of $\mathcal{D}$. If $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2},\epsilon,\frac{\delta}{2})$-stable.
@@
We may assume that $\frac{1}{n}\leq\frac{1}{n_{0}}$ where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is the constant from Lemma 2.2. Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $[n]$ and a digraph collection $\mathcal{D}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for each $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq n\leq Mm+\epsilon_{0}n$ and $Mm\leq n\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Given $X\subset[L]$, we write $\hat{X}:=\bigcup_{j\in X}V_{j}$. Hence $|\hat{X}|=m|X|$.
@@
Let $d\ll\alpha^{\prime}$ and $i\in[M]$. Assume that there are sets $A,B\subset[L]$ of size at least $(\frac{1}{2}-\alpha^{\prime})L$ such that $e_{R_{i}}(A,B)\leq\alpha^{\prime}L^{2}$. Then $e_{D_{c}}(\hat{A},\hat{B})\leq\sqrt{\alpha^{\prime}}n^{2}$ for all but at most $2\sqrt{\alpha^{\prime}}m$ colors $c\in\mathcal{C}_{i}$.
@@
Let $t$ be the number of colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A},\hat{B})>\sqrt{\alpha^{\prime}}n^{2}$. Since $e(D_{c})-e(D_{c}^{\prime})\leq(24d+\epsilon_{0})n^{2}$ for all $c\in[n]$, we have
$\sqrt{\alpha^{\prime}}n^{2}t$ $\leq\sum_{c\in\mathcal{C}_{i}}e_{D_{c}}(\hat{A},\hat{B})\leq\sum_{c\in\mathcal{C}_{i}}\left(e_{D_{c}^{\prime}}(\hat{A},\hat{B})+(24d+\epsilon_{0})n^{2}\right)$
$\leq|\mathcal{C}_{i}|e_{R_{i}}(A,B)m^{2}+|\mathcal{C}_{i}|(24d+\epsilon_{0})n^{2}$
$\leq\alpha^{\prime}L^{2}m^{3}+25dmn^{2}\leq 2\alpha^{\prime}mn^{2}.$
@@
We consider the cases where $\mathcal{D}$ is strongly stable and weakly stable, respectively.
@@
We prove that $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Suppose that there exists a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$. By Definition 3.1, for each $i\in I$, there are two sets $A^{i},B^{i}\subseteq[L]$ of size at least $(\frac{1}{2}-\alpha^{2})L$ such that $e_{R_{i}}(A^{i},B^{i})<\alpha^{2}L^{2}$. Applying Claim 4.2 with $\alpha^{\prime}:=\alpha^{2}$, we get that there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A^{i}},\hat{B^{i}})\leq\alpha n^{2}$. Notice that $|\hat{A^{i}}|,|\hat{B^{i}}|\geq(\frac{1}{2}-\alpha^{2})Lm>(\frac{1}{2}-\alpha)n$. Then there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i}$ such that $D_{c}$ is not $\alpha$-nice. Thus the number of colors $c\in[n]$ for which $D_{c}$ is not $\alpha$-nice is at least $(1-2\alpha)m|I|>(1-\gamma)n$ since $\alpha\ll\gamma$. This contradicts the fact that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable.
@@
Suppose that $\mathcal{R}$ is not $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Then there is a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$. It suffices to show that $\mathcal{R}$ is $(\epsilon,\frac{\delta}{2})$-weakly stable. Choose $i\in I$, then $R_{i}$ is $\alpha^{2}$-extremal. Based on Lemma 2.4 (i), one has $d_{R_{i}}^{+}(j),d_{R_{i}}^{-}(j)\geq(\frac{1}{2}-2\mu)L$ for all but at most $d^{\frac{1}{4}}L$ vertices $j\in[L]$. Together with Lemma 3.2, we obtain that
  • if $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, then $R_{i}$ has a characteristic partition $(A_{i}^{\prime},B_{i}^{\prime},L_{i}^{\prime})$ such that there are $Z,Y\in\{A,B\}$ satisfying $e_{R_{i}}(Z_{i}^{\prime},Y_{i}^{\prime})\leq\alpha^{2}L^{2}$;
  • if $R_{i}$ is $(\alpha^{2},{\rm EC}3)$-extremal, then $R_{i}$ has a characteristic partition $({C_{1}^{i}}^{\prime},{C_{2}^{i}}^{\prime},{C_{3}^{i}}^{\prime},{C_{4}^{i}}^{\prime},L_{i}^{\prime})$ such that $e_{R_{i}}(Z_{i}^{\prime},Y_{i}^{\prime})\leq\alpha^{2}L^{2}$ with $(Z_{i}^{\prime},Y_{i}^{\prime})\in\{({W_{i}^{4}}^{\prime},{W_{i}^{3}}^{\prime}),({W_{i}^{2}}^{\prime},{W_{i}^{1}}^{\prime})\}$.
Note that $|\hat{Z}_{i}^{\prime}|,|\hat{Y}_{i}^{\prime}|=(\frac{1}{2}-\alpha^{2})Lm\geq(\frac{1}{2}-2\alpha)n$. In view of Claim 4.2, we know that there is a subset $\mathcal{B}_{i}\subseteq\mathcal{C}_{i}$ with $|\mathcal{C}_{i}\setminus\mathcal{B}_{i}|\leq 2\alpha m$ such that for all $c\in\mathcal{B}_{i}$, we have $e_{D_{c}}(\hat{Z}_{i}^{\prime},\hat{Y}_{i}^{\prime})\leq\alpha n^{2}$. It follows that $D_{c}$ is $2\alpha$-extremal for each $c\in\mathcal{B}_{i}$.
If $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, then each $D_{c}$ with $c\in\mathcal{B}_{i}$ admits a characteristic partition $(A_{c},B_{c},L_{c})$, and there are $W\in\{A,B\}$ and $\{Z,Y\}=\{A,B\}$ such that $e_{D_{c}}(Z_{c},W_{c})\leq\sqrt{\alpha}n^{2}$ and $e_{D_{c}}(Y_{c},W_{c})\geq(\frac{1}{4}-\sqrt{\alpha})n^{2}$. Hence, we must have either $|\hat{Z}_{i}^{\prime}\triangle Z_{c}|,|\hat{Y}_{i}^{\prime}\triangle Y_{c}|\leq\epsilon n$ or $|\hat{Z}_{i}^{\prime}\triangle Y_{c}|,|\hat{Y}_{i}^{\prime}\triangle Z_{c}|\leq\epsilon n$. So, either $|\hat{A}_{i}^{\prime}\triangle A_{c}|,|\hat{B}_{i}^{\prime}\triangle B_{c}|\leq\epsilon n$ or $|\hat{A}_{i}^{\prime}\triangle B_{c}|,|\hat{B}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$. That is, for all $c\in\mathcal{B}_{i}$, the characteristic partition of $D_{c}$ is almost the same as the union of clusters corresponding to the characteristic partition of $R_{i}$.
If $R_{i}$ is $(\alpha^{2},{\rm EC}3)$-extremal, then each $D_{c}$ admits a characteristic partition $(C_{c}^{1},C_{c}^{2},C_{c}^{3},C_{c}^{4},L_{c})$ for $c\in\mathcal{B}_{i}$, and it satisfies $e_{D_{c}}(W_{c}^{2},W_{c}^{1}),e_{D_{c}}(W_{c}^{4},W_{c}^{3})\leq\sqrt{\alpha}n^{2}$, $e_{D_{c}}(C_{c}^{1}),e_{D_{c}}(C_{c}^{3})\geq|C_{c}^{1}|^{2}-\sqrt{\alpha}n^{2}$ and $e_{D_{c}}(C_{c}^{2},C_{c}^{4}),\,e_{D_{c}}(C_{c}^{4},C_{c}^{2})\geq|C_{c}^{2}|^{2}-\sqrt{\alpha}n^{2}$. Therefore, either $|\hat{Z}_{i}^{\prime}\triangle W_{c}^{2}|,|\hat{Y}_{i}^{\prime}\triangle W_{c}^{1}|\leq\epsilon n$ or $|\hat{Z}_{i}^{\prime}\triangle W_{c}^{4}|,|\hat{Y}_{i}^{\prime}\triangle W_{c}^{3}|\leq\epsilon n$. By relabeling, one may assume that $|{\hat{W_{i}^{4}}^{\prime}}\triangle W_{c}^{4}|,|{\hat{W_{i}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$ for each $c\in\mathcal{B}_{i}$. That is, for all $c\in\mathcal{B}_{i}$, the characteristic partition of $D_{c}$ is almost the same as the union of clusters corresponding to the characteristic partition of $R_{i}$.
Choose $c\in\mathcal{B}_{i},c^{\prime}\in\mathcal{B}_{j}$ for some $i,j\in[L]$. Suppose that $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing. We must have $i\neq j$. We consider the following possible cases.
$\bullet$ Suppose that $R_{i}$ and $R_{j}$ are either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal. If $|\hat{A}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$ and $|\hat{A}_{j}^{\prime}\triangle A_{c^{\prime}}|\leq\epsilon n$, then
$|A_{i}^{\prime}\triangle A_{j}^{\prime}|m=|\hat{A}_{i}^{\prime}\triangle\hat{A}_{j}^{\prime}|\geq|A_{c}\triangle A_{c^{\prime}}|-|\hat{A}_{i}^{\prime}\triangle A_{c}|-|\hat{A}_{j}^{\prime}\triangle A_{c^{\prime}}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
@@
$\bullet$ Suppose that $R_{i}$ and $R_{j}$ are $(\alpha^{2},{\rm EC}3)$-extremal. Recall that $|{\hat{W_{i}^{4}}^{\prime}}\triangle W_{c}^{4}|,|{\hat{W_{i}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$. Then $|\hat{W_{i}^{1}}^{\prime}\triangle W_{c}^{1}|\leq 10\epsilon n$ and $|\hat{W_{i}^{2}}^{\prime}\triangle W_{c}^{2}|\leq 10\epsilon n$. Since $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing, one may assume, without loss of generality, that $|W_{c}^{1}\triangle W_{c^{\prime}}^{1}|\geq\delta n$ and $|W_{c}^{1}\triangle W_{c^{\prime}}^{3}|\geq\delta n$. Hence
$|{W_{i}^{1}}^{\prime}\triangle{W_{j}^{1}}^{\prime}|m=|{\hat{W_{i}^{1}}^{\prime}}\triangle{\hat{W_{j}^{1}}^{\prime}}|\geq|W_{c}^{1}\triangle W_{c^{\prime}}^{1}|-|{\hat{W_{i}^{1}}^{\prime}}\triangle W_{c}^{1}|-|{\hat{W_{j}^{1}}^{\prime}}\triangle W_{c^{\prime}}^{1}|\geq\delta n-21\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
and similarly, $|{W_{i}^{1}}^{\prime}\triangle{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
$\bullet$ Suppose that $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, and $R_{j}$ is $(\alpha^{2},{\rm EC}3)$-extremal. If $|\hat{A}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$ and $|\hat{{W_{j}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$, then
$|A_{i}^{\prime}\triangle{W_{j}^{3}}^{\prime}|m=|\hat{A}_{i}^{\prime}\triangle\hat{{W_{j}^{3}}^{\prime}}|\geq|A_{c}\triangle W_{c^{\prime}}^{3}|-|\hat{A}_{i}^{\prime}\triangle A_{c}|-|\hat{{W_{j}^{3}}^{\prime}}\triangle W_{c^{\prime}}^{3}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
Similarly, $|B_{i}^{\prime}\triangle{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
@@
For a digraph collection $\mathcal{D}$ that is either strongly stable or weakly stable, Lemma 3.6 and Lemma 3.9 guarantee the existence of a directed absorbing cycle in $\mathcal{D}$. We shall then employ this absorbing property to prove that $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
Let $0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. If $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
Choose additional parameters $n_{0},L_{0},\epsilon_{0},d,\beta,\lambda$, where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is obtained from Lemma 2.2, such that
$1<\frac{1}{n}\ll\frac{1}{n_{0}}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu\ll\beta\ll\lambda\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1.$
@@
For this, we first count the number of choices for $v_{1}$. Let $N_{1}:=\{x\in N_{{c}}^{-}(u_{1}):x\ \text{is}\ D_{b_{1}}\text{-good}\}$. Hence, $|N_{1}|\geq\left(\frac{1}{2}-\mu\right)n-2\epsilon n$. By (b), we know that all but at most $\sqrt{\epsilon}n$ vertices $v_{1}\in N_{1}$ satisfying that $(b_{1},v_{1})$ is absorbable. Hence there are at least $\frac{n}{4}$ choices of $v_{1}$ such that $v_{1}\in N_{1}$ and $(b_{1},v_{1})$ is absorbable.
Now fix $v_{1}$. Let $N_{2}:=\{x\in N_{{b_{2}}}^{+}(u_{2}):x\ \text{is}\ D_{b_{2}}\text{-good}\}$. Hence, $|N_{2}|\geq\left(\frac{1}{2}-\mu\right)n-2\epsilon n$. By (a) and (b), we know that all but at most $2\sqrt{\epsilon}n+1$ vertices $v_{2}\in N_{2}$ such that $v_{2}\neq v_{1}$ and $(b_{1},b_{2},v_{1},v_{2})$ is totally absorbable. Therefore, $|S(b_{1},b_{2},u_{1},u_{2},c)|\geq 2^{-4}n^{2}$, as desired.
Let $\{c_{i},c_{i}^{\prime}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a collection of distinct colors. For each $i\in[2\beta n]$, let $\mathcal{F}_{i}$ be the directed 2-graph on $V$ with edge set $\{(v_{1},v_{2})\in V^{2}:(c_{i},c_{i}^{\prime},v_{1},v_{2})\ \text{is totally absorbable}\}$, and define $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$. Define the collection of directed multi-graphs:
@@
For every $S:=S(u_{1},u_{2},c)\in\mathbf{Z}$ and every $i\in[2\beta n]$, we have $|E(S)\cap E(\mathcal{F}_{i})|\geq 2^{-4}n^{2}$. Applying Lemma 3.4 with $t:=2\beta n$ and $\epsilon:=2^{-4}$, we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-9}\beta n$ for all $S\in\mathbf{Z}$, as desired. ∎